2w + 5w +6w -3w +5w -4w= 70..........w=?
Found 2 solutions by checkley79, Tatiana_Stebko:
Answer by checkley79(3341)

2w + 5w +6w -3w +5w -4w= 70
11w=70
w=70/11=6.3636 ans.

Answer by Tatiana_Stebko(1539)